OpenStudy (shesolitt):

Write the equation of the line that passes through (1, 5) and (–2, 14) in slope-intercept form. i will medal and fan !

pooja195 (pooja195):

First find the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (shesolitt):

y = 3x + 2 y = 3x + 8 y = –3x – 2 y = –3x + 8

pooja195 (pooja195):

Find the slope =)

OpenStudy (shesolitt):

i have to plug it in ? @pooja195

pooja195 (pooja195):

\[\LARGE \frac{14-5= ? }{-2-1=?}=slope\]

OpenStudy (anonymous):

wut

OpenStudy (anonymous):

hi @pooja195

pooja195 (pooja195):

@shesolitt have you figured it out? And Hi @Gabriel_Morrow

OpenStudy (shesolitt):

so 9 and -1

pooja195 (pooja195):

-2-1=? its not -1

OpenStudy (shesolitt):

i meant -3

pooja195 (pooja195):

ok 9/-3=?

OpenStudy (shesolitt):

-3

OpenStudy (anonymous):

...

OpenStudy (shesolitt):

@pooja195

pooja195 (pooja195):

Yes i need to go perhaps @jabez177 or @imqwerty can take over hopefully sorry this is the end of my free period!

OpenStudy (anonymous):

@imqwerty @jabez177 help

imqwerty (imqwerty):

if u have 2 points (x1,y1) and (x2,y2) then the equation of line is given like-\[\frac{ y-y_{1} }{ y_{2}-y_{1} }=\frac{ x-x_{1} }{ x_{2}-x_{1} }\] 1st u find the equation by puting values in this^ then convert it to this form-\[y=mx+c\]

OpenStudy (shesolitt):

so i have to plug them in all over again

jabez177 (jabez177):

Yes. Or just read it over and finish and continue plugging it in from where Pooja left off.

OpenStudy (shesolitt):

im confused where to plug it in

OpenStudy (shesolitt):

i got 9 and -3

imqwerty (imqwerty):

you have (x1,y1)=(1, 5) and (x2,y2)= (–2, 14) plug them in and follow the steps

OpenStudy (shesolitt):

9 and -3 ?....

pooja195 (pooja195):

Wow this is plain trolling they had the correct answer im so sorry @shesolitt I expected better from these people yes once you divide your slope is -3 Now you can convert into y=mx+b form.